OpenStudy (anonymous):

What is the range of the graph of y = -7(x - 1)2 + 2? y greater than or equal to 7 y less than or equal to 7 y less than or equal to 2 y greater than or equal to 2

OpenStudy (aum):

What is the vertex of this graph (it is a parabola)?

OpenStudy (anonymous):

no cluuueeee

OpenStudy (anonymous):

first simplify

OpenStudy (aum):

If y = a(x-h)^2 + k then (h,k) is the vertex.

OpenStudy (anonymous):

open bracket

OpenStudy (anonymous):

simplify nothing

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[y = -7(x - 1)^2 + 2\]think only \((x-1)^2\geq 0\) because it is a square \[-7(x-1)^2\leq 0\] because of the \(-7\) out front so \[-7(x-1)^2+2\leq 2\]

OpenStudy (anonymous):

so which one is the answer

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@driftracer305

OpenStudy (driftracer305):

the max value for y here is 2

OpenStudy (anonymous):

So is it C or D?

OpenStudy (driftracer305):

so if the max is 2........ can something go higher than a MAX?

OpenStudy (anonymous):

ok its C, THANK YOU

OpenStudy (driftracer305):

hey...... u got it on ur own........ so great work....